Cookies on trays problem (by M Pollino) Exactly 7 cookies fit on each tray for baking. Four-sevenths of a tray is used to make a small box of cookies. If a baker has made 8 trays of cookies, how many boxes can be filled?


1
Expert's answer
2021-06-22T17:42:56-0400

No. Of cookies fit in each tray= 7

four - sevenths of a tray is used to make a small box of cookies.

Total number of trays= 8

According to given,

Number of cookies in a box= 7×(4÷7)= 4 cookies

Total number of cookies made= Number of trays* Number of cookies in each tray= 8×7= 56

Therefore,

Total number of boxes= Total number of cookies/Number of cookies in each box

= 56÷4= 14

Hence,

A total of 14 boxes can be filled.